# A Calculus of Overlays

### v2

@@ 6.1. Emergent Phenomena
Each of these features is primitive in overlay-calculus. In the $\lambda$-calculus, encoding them requires pervasive changes to program structure: every scope must thread a mutable store for lazy allocation, every method invocation must consult a dynamically constructed dictionary for open recursion, and every composition must be wrapped in order-insensitive infrastructure for commutativity. These are global, cascading transformations, not local rewrites of individual terms.
6.2. Classification of Semantic Variants
The goals of declarative programming rule out several properties of the model: separate primitives for methods, classes, types, and objects; mutable reference cells; domain-theoretic approximation; and asymmetric override. Ruling these out leaves a minimal syntax of three constructs: records, definitions, and inheritance.
Two degrees of freedom remain. These dimensions are independent, and each combination yields a coherent semantics. They arise in any system whose definitions form a tree of named scopes, where each node is a path and references are path-dependent, even in systems ruled out above, such as Scala, which has separate primitives for methods, classes, objects, and types.
Existing designs occupy several of the resulting cells:
No deep merge Deep merge
Early binding CUE (van Lohuizen, 2019)
Dynamic scope union file systems, Jsonnet (Cunningham, 2014)
Lexical scope Scala (Odersky et al., 2004) NixOS modules (NixOS Contributors, 2024)
Multi-target $\mathtt{this}$ overlay-calculus
Early binding
With early binding, references resolve at definition time to a fixed path, independently of how the record is later inherited. CUE (van Lohuizen, 2019) exemplifies this: field references are statically resolved, and a field in a nested record cannot refer back to properties of its enclosing context contributed by later composition. Early binding combined with deep merge is coherent and practically useful, but cross-overlay cooperation requires explicit wiring rather than inheritance: a definition cannot observe properties contributed by overlays composed after it.
Dynamic scope
With dynamic scope, references resolve against the context at the point of use rather than the definition site. Early Lisp implementations used dynamic scope (McCarthy, 1978), where variable references are resolved against the run-time call stack. Union file systems exhibit the same property at the file-system level: a symbolic link’s relative path resolves against the mount context at dereference time, not against the layer that defined the link. Jsonnet (Cunningham, 2014) follows this approach: self always refers to the final merged object, not the object at the definition site. When the Scala multi-path example (Appendix B) is translated to Jsonnet, it does not error, but outer resolves to the merged MyObjectA rather than to Object1 or Object2 separately, silently discarding the distinction between the two inheritance routes. Dynamic scope is flexible but loses the ability to track which inheritance route a reference came from.
Deep merge
Without the second case of equation (3), $\mathtt{overlays}(p)$ would always return $\{p\}$, a single element. When two independently authored overlays both define a label with the same name at the same scope level, only the definition at the literal path $p$ would contribute to $\mathtt{supers}$ and hence to $\mathtt{properties}$; the other definitions would be silently invisible. This is the situation in most composition systems: pytest permits at most one fixture per name per scope; trait systems (Schärli et al., 2003) flag same-name methods as conflicts. Under any of these rules, two independently authored overlays that both extend an entity with the same name, which is the situation that arises naturally when solving the Expression Problem (Wadler, 1998), cannot be composed. The non-trivial case of equation (3) removes this limitation: it collects all paths that share the same label and merges their definitions rather than shadowing each other. This is the record-level counterpart of method override in object-oriented languages, and the mechanism by which the calculus admits extensibility in both dimensions of the Expression Problem (Wadler, 1998) (Section 5).
Multi-target qualified this
The qualified this reference $\mathtt{this}$ answers the question: “in the fully inherited tree, where does the definition-site scope $p_{\mathrm{def}}$ actually live?” Scheme (Sussman and Steele Jr., 1975) introduced lexical scope, where a de Bruijn index resolves in a single step by indexing into the statically determined environment; this is predictable but rigid. Scala (Odersky et al., 2004) and the NixOS module system (NixOS Contributors, 2024) follow the lexical approach: each this reference is bound to a single statically known path. When inheritance introduces multiple inheritance routes to the same scope, Scala rejects the ambiguity at compile time (Appendix B) and the NixOS module system rejects the duplicate option declaration (Appendix B); neither can express that all routes are equally valid.
Overlay-calculus combines the predictability of lexical scope with the flexibility of dynamic scope. The reference itself is lexically determined at the definition site as a de Bruijn index, so what $\mathtt{this}$ looks up is predictable. But because inheritance interleaves scopes from different overlays, the result depends on the inherited tree structure: $\mathtt{this}$ walks upward one scope level at a time, consulting $\mathtt{supers}$ at each level, and naturally tracks a set of inheritance-site paths rather than a single one. All routes contribute equally to the inherited result, and since inheritance is idempotent, duplicated contributions are harmless. The only observable is property presence, making multi-path resolution well-defined without disambiguation.
@@ 6.2. Classification of Semantic Variants
Our first three implementations of the Overlay language used standard techniques for self-reference resolution: closure-based fixed points following Cook (Cook, 1989), and stack-based environment lookup using de Bruijn indices (de Bruijn, 1972). All three produced subtle bugs when inheritance introduced multiple routes to the same enclosing scope. The difficulty was structural, not incidental: a closure captures a single environment, so a fixed-point combinator solves for a single self-reference; a stack-based environment is a linear chain where each scope level has one parent. Both data structures embed a single-valued assumption incompatible with the multi-path situation that multiple inheritance naturally produces (Appendix B). Equation (6) emerged from abandoning this assumption: it tracks a frontier set $S$ rather than a single current scope. If the equations were restricted to single-target resolution, as in DOT’s single-valued self variable (Amin et al., 2016), they would need to reject structurally legitimate cases, adding complexity rather than removing it.
Consequences of the alternative choices
Choosing differently in any dimension requires introducing machinery that is absent from the equations of Section 3. With early binding, the equations must freeze reference targets at definition time, so a definition cannot observe properties contributed by later-composed overlays. With dynamic scope, the equations must resolve references against a runtime context rather than the definition-site structure, losing the ability to distinguish inheritance routes. Restricting $\mathtt{overlays}(p)$ to $\{p\}$ means same-label definitions from different inheritance sources no longer share identity, requiring either a method resolution order (MRO) (Barrett et al., 1996) to linearize them or a shadowing rule to select one and discard the rest, which is why systems without deep merge rarely support path-dependent multiple inheritance in practice. Restricting $\mathtt{this}$ to a single path requires a mechanism to select among inheritance routes (Scala rejects the ambiguity at compile time; the NixOS module system raises a duplicate-declaration error), and loses contributions from all non-selected routes. These mechanisms, including early-binding freezing, runtime context threading, MRO, shadowing, route selection, and error reporting, are more artificial than their absence. The equations of Section 3 contain none of them.
@@ 7. Related Work
Amin et al. (Amin et al., 2016) formalized path-dependent types in the DOT calculus, the theoretical foundation of Scala’s type system. DOT paths navigate through objects to reach nested type members, providing a rich dependent type structure. DOT’s self variable in the object constructor $\nu(x:T)\,d$ binds $x$ to a single object; when multiple inheritance routes lead to the same enclosing scope, the resulting intersection type is unrealizable and Scala rejects it at compile time (Appendix B). DOT stratifies its syntax into terms and types, with types further divided into abstract type members bounded by $L..U$ and concrete definitions; each stratum requires its own subtyping rules. Whether these three roles could be unified under a single construct remains an open question in the DOT line of work. DOT includes the $\lambda$-calculus as a sub-language; removing it while retaining path-dependent types would yield a calculus closer to pure overlay inheritance (see Section 8).
@@
The NixOS module system (Dolstra et al., 2010; NixOS Contributors, 2024) composes modules by recursively merging nested attribute sets, not by linearization or method-level override, which is the same overlay semantics that distinguishes tree-level inheritance from mixin and trait composition. When two modules define the same nested path, their subtrees are merged rather than overridden. The NixOS module system incorporates deep merge, recursive self-reference, deferred evaluation of modules, and a rich type-checking layer on top. The same mechanism powers NixOS system configuration, Home Manager (Helgesson, 2017), nix-darwin (Jordan, 2017), flake-parts (Hensing, 2022), and dozens of other ecosystems, collectively managing configurations of arbitrary complexity. The one aspect that appears not to have been fully addressed is the semantics of self-reference under multiple inheritance. When two independent modules that inherit from the same deferred module are composed, each import introduces its own declaration of the shared options; the module system’s mergeOptionDecls rejects duplicate declarations with a static error (Appendix B). The module system cannot express that both routes contribute equally to the same option, that is, the multi-path situation that recursive merging naturally gives rise to, which motivated the coalgebraic semantics of Section 3.
@@
Plugin systems, component frameworks, and dependency injection frameworks enable software extensibility through declarative inheritance. The symmetric, associative nature of inheritance (Section 3) explains why such systems work: components can be inherited in any order without changing behavior. These systems implement overlay-calculus patterns ad hoc.
@@
Union file systems (UnionFS, OverlayFS, AUFS) layer multiple directory hierarchies to present a unified view. Their semantics resembles overlay-calculus in that later layers override earlier layers and files from all layers remain accessible, but without a formal foundation these systems resort to ad-hoc choices: asymmetric inheritance, non-commutative conflict resolution for scalar files, and no late-binding or dynamic dispatch across layers. We implemented overlay-calculus directly as a union file system (Yang, 2025a), obtaining late-binding semantics and dynamic dispatch for file lookups with no additional mechanism beyond the three primitives. In package managers, build systems, and OS distributions, external toolchains transform configuration into file trees; here, the file system serves as configuration to itself, and the transformation is inheritance: a file-system-as-compiler in which the directory tree is both the source program and the compilation target.

### v4

@@ 6.1. Emergent Phenomena
Each of these features is primitive in inheritance-calculus. In the $\lambda$-calculus, encoding them requires pervasive changes to program structure: every scope must thread a mutable store for lazy allocation, every method invocation must consult a dynamically constructed dictionary for open recursion, and every composition must be wrapped in order-insensitive infrastructure for commutativity. These are global, cascading transformations, not local rewrites of individual terms. In common sense, inheritance-calculus is strictly more expressive than the $\lambda$-calculus.
6.2. Semantic Alternatives
The preceding item established that inheritance-calculus is immune to nonextensibility. We now ask whether an alternative design could share this property by surveying every candidate known to us.
Degrees of freedom
Every language whose programs are tree-structured ASTs with references can be characterized by two finite sets: a set $\mathcal{N}$ of node types, each determining how a subtree organizes its children, and a set $\mathcal{R}$ of reference types, each determining how one position in the tree refers to another. The syntax is fixed by the choice of $\mathcal{N}$ and $\mathcal{R}$; the semantics is determined by how each node type structures composition and how each reference type resolves its target.
Language Node types $\mathcal{N}$ Reference types $\mathcal{R}$
Java (OOP) class, interface, method, block, conditional, loop, throw, lambda, … variable, field access, method call, extends/implements, import, …
Haskell (FP) $\lambda$, application, let, case, data, type class, instance, … variable, constructor, type class dispatch, import, …
Scala (multi-paradigm) class, trait, object, method, block, conditional, match, while, throw, … variable, member access, extends, with, implicits, import, …
$\lambda$-calculus abstraction, application variable
Inheritance-calculus record inheritance
Immunity to nonextensibility
A path $p$ in a program $P$ is extensible if there exists a program $Q$ such that composing $P$ with $Q$ through a reference makes new properties observable at $p$ without modifying $P$. A language is immune to nonextensibility if every path in every well-formed program is extensible.
We now classify the design space to explain why we chose the semantics of Section 3. Two axes remain free: the composition mechanism—how the single reference type incorporates definitions; and the reference resolution—how the single reference type finds its target.
The composition mechanism
Every composition mechanism determines how the definitions of one abstraction are incorporated into another. The following table classifies the known candidates by whether they satisfy commutativity (C), idempotence (I), and associativity (A), and by their extensibility model.
Candidate C I A Extensibility Representative
Deep merge $\checkmark$ $\checkmark$ $\checkmark$ universal this paper
Shallow merge $\times$ $\checkmark$ $\checkmark$ top-level only JS spread
Conflict rejection $\checkmark$ $\checkmark$ $\checkmark$ none on overlap Harper–Pierce (Harper and Pierce, 1991)
Priority merge $\times$ $\times$ $\checkmark$ universal Jsonnet +, Nix //
$\beta$-reduction $\times$ $\times$ $\times$ opt-in Java, Haskell, Scala, …
Conditional breaks monotonicity
Deep merge:
Merging same-label definitions recurses into their subtrees ($\mathtt{overrides}$, equation 3). Every label is an extension point, whether or not the original author anticipated extension.
Shallow merge:
Nested labels cannot be independently extended.
Conflict rejection:
Same-label definitions are rejected as errors. All three identities hold—but vacuously: the identities are satisfied because the compositions that would test them are refused. Conflict rejection prevents extending existing labels (Harper and Pierce, 1991)—the composability that the Expression Problem (Wadler, 1998) requires.
Priority merge:
$A+B\neq B+A$: one source takes precedence (Cunningham, 2014; Dolstra et al., 2010), introducing an ordering dependency.
$\beta$-reduction:
Only the parameters placed by the author are extension points; the function body is closed. To make a new aspect extensible, the original function must be rewritten to accept an additional parameter. This applies to every language whose composition mechanism is function application—functional, object-oriented, and multi-paradigm alike—because method calls, virtual dispatch, and type-class resolution all reduce to $\beta$-reduction. The Expression Problem (Wadler, 1998) is hard in all of them for the same reason: extensibility is opt-in. The frameworks that simulate universal extensibility (visitor pattern, object algebras (d. S. Oliveira and Cook, 2012), finally tagless interpreters (Carette et al., 2009)) are the machinery needed to work around this limitation.
Conditional incorporation:
Removing or conditionally including definitions breaks monotonicity, so the semantic equations (Appendix C) may have no fixed point (Tarski, 1955).
Among the candidates above, deep merge is the only one we found that is commutative, idempotent, associative, and universally extensible. The other candidates each introduce a vulnerability to nonextensibility or break monotonicity.
The reference resolution
Once deep merge over records is fixed as the composition mechanism, four of the six semantic equations are determined by the syntax: $\mathtt{properties}$ (1) reads the labels of a record, $\mathtt{bases}$ (4) reads its inheritance sources, $\mathtt{supers}$ (2) takes the transitive closure, and $\mathtt{overrides}$ (3) implements the merge. The only remaining degree of freedom is how references resolve: the $\mathtt{resolve}$ (5) and $\mathtt{this}$ (6) equations. We examine the three known candidates.
Early binding:
With early binding, references resolve at definition time to a fixed path, independently of how the record is later inherited. CUE (van Lohuizen, 2019) exemplifies this: field references are statically resolved, and a nested field cannot refer to properties of its enclosing context contributed by later composition. This prevents the mixin from serving as a call stack for the $\lambda$-calculus: $\beta$-reduction requires the callee to observe the argument supplied at the call site, which is a later composition. With early binding, references are frozen before this composition occurs. In the equations, $\mathtt{resolve}$ (5) calls $\mathtt{this}$ (6), which walks the inherited tree structure; late binding is what allows $\mathtt{this}$ to see properties contributed by later-composed mixins. Without a Turing-complete mixin, functions must be reintroduced as the composition mechanism, re-exposing the opt-in extensibility problem above.
Dynamic scope:
With dynamic scope, references resolve at the point of use rather than the definition site. Jsonnet (Cunningham, 2014) follows this approach: self always refers to the final merged object, not the object at the definition site. Union file systems exhibit the same property at the file-system level: a symbolic link’s relative path resolves against the mount context at dereference time, not against the layer that defined the link. Dynamic scope breaks the $\lambda$-calculus embedding (Section 4). The Substitution Lemma (Lemma D.2) relies on the base case $M=y$ ($y\neq x$): the de Bruijn index $m$ of $y$ walks up from the definition-site path $\mathtt{init}(p_{\mathrm{def}})$ to a scope level different from $x$’s binder, so the substitution $\mathrm{argument}\mapsto\mathcal{T}(V)$ at $x$’s scope has no effect. With dynamic scope this separation fails. [^4] Counterexample: $(\lambda x.\,\lambda y.\,x)\;V_{1}\;V_{2}$. The reference to $x$ inside $\lambda y.\,x$ translates to $\uparrow^{\!1}.\mathrm{argument}$ (de Bruijn index $1$). With late binding, $\mathtt{this}$ starts from the definition-site scope (the inner $\lambda$) and walks up one level to the outer $\lambda$, reaching the slot $\mathrm{argument}\mapsto\mathcal{T}(V_{1})$. With dynamic scope, the reference resolves from the use-site context, where the innermost enclosing scope is the application $\{C_{1}.\mathrm{result},\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\}$; walking up one level reaches this scope’s $\mathrm{argument}$ slot, returning $\mathcal{T}(V_{2})$ instead of $\mathcal{T}(V_{1})$. This is the classical variable-capture failure of dynamic scoping. As with early binding, the mixin cannot serve as a complete replacement for functions.
Single-target $\mathtt{this}$:
Single-target $\mathtt{this}$ is a promising candidate: the Single-Path Lemma (D.1) shows that for translated $\lambda$-terms, the frontier set $S$ in $\mathtt{this}$ (6) always contains exactly one path, so the $\lambda$-calculus embedding does not require multi-target resolution. However, single-target $\mathtt{this}$ is vulnerable to nonextensibility for a different reason. When multiple inheritance routes reach the same scope, single-target resolution must reject the situation or select one route. The NixOS module system (NixOS Contributors, 2024) raises a duplicate-declaration error (Appendix B). This rejects nested extensions where two independently authored modules define the same option—a composition that arises when solving the Expression Problem (Wadler, 1998) (Section 5).
@@ 6.2. Semantic Alternatives
Our first three implementations of MIXINv2 used standard techniques for self-reference resolution: closure-based fixed points following Cook (Cook, 1989), and stack-based environment lookup using de Bruijn indices (de Bruijn, 1972). All three produced subtle bugs when inheritance introduced multiple routes to the same enclosing scope. The difficulty was structural: a closure captures a single environment, so a fixed-point combinator solves for a single self-reference; a stack-based environment is a linear chain where each scope level has one parent. Both data structures embed a single-valued assumption incompatible with the multi-path situation that multiple inheritance naturally produces (Appendix B). Equation (6) emerged from abandoning this assumption: it tracks a frontier set $S$ rather than a single current scope.
The following table summarizes the vulnerabilities of all alternatives at both levels.
Candidate Vulnerability Representative
composition Shallow merge nested extensions not composable JS spread
Conflict rejection same-label extensions rejected Harper–Pierce (Harper and Pierce, 1991)
Priority merge ordering dependency Jsonnet + (Cunningham, 2014), Nix // (Dolstra et al., 2010)
$\beta$-reduction opt-in extensibility Java, Haskell, Scala, …
Conditional breaks monotonicity
resolution Early binding mixin cannot replace function CUE (van Lohuizen, 2019)
Dynamic scope mixin cannot replace function Jsonnet (Cunningham, 2014), union FS
Single-target $\mathtt{this}$ nested extensions rejected NixOS modules (NixOS Contributors, 2024)
No alternative we examined at either level is immune to nonextensibility.
The semantics of inheritance-calculus was not chosen from a menu. We arrived at it by trying each alternative in this section and finding that it failed: shallow merge could not compose nested extensions; priority merge introduced ordering dependencies; early binding froze references before composition could supply them; dynamic scope broke the substitution that the $\lambda$-calculus embedding relies on; single-target $\mathtt{this}$ rejected the multi-path compositions that the Expression Problem demands (three successive implementations confirmed this—see above). Deep merge with multi-target late-binding $\mathtt{this}$ is what remained.
@@ 7. Related Work
Semantic descriptions of the $\lambda$-calculus
The semantics of the $\lambda$-calculus can be defined in several fundamentally different ways. Operational semantics takes $\beta$-reduction $(\lambda x.\,M)\,N\to M[N/x]$ as the primitive notion of computation (Church, 1936, 1941; Barendregt, 1984). This requires capture-avoiding substitution, a mechanism whose subtleties motivated de Bruijn indices (de Bruijn, 1972) and explicit substitution calculi (Abadi et al., 1991), and a choice of evaluation strategy: call-by-name versus call-by-value (Plotkin, 1975) or the lazy strategy of Abramsky (Abramsky, 1990). Denotational semantics (Scott, 1970, 1976) interprets terms in a mathematical domain, but giving meaning to $\beta$-reduction requires a reflexive domain $D\cong[D\to D]$, which has no set-theoretic solution by Cantor’s theorem; Scott’s $D_{\infty}$ resolves this by constructing $D$ as an inverse limit of continuous lattices (Scott, 1972). The resulting semantics is adequate but not fully abstract (Plotkin, 1977; Milner, 1977): it identifies terms that observational equivalence distinguishes (Plotkin, 1975). Game semantics (Abramsky et al., 2000; Hyland and Ong, 2000; Nickau, 1994) achieves full abstraction for PCF and for the lazy $\lambda$-calculus (Abramsky and Ong, 1993; Abramsky and McCusker, 1995) at the cost of considerably more intricate mathematical machinery (arenas, strategies, innocence conditions). In each of these approaches, the function space $D\to D]$ is the central obstacle: operational semantics needs it implicitly (via substitution), denotational semantics constructs it explicitly, and game semantics refines it via strategies. None of these approaches yields a semantics of the lazy $\lambda$-calculus that simultaneously lives in naive set theory and is fully abstract: Scott’s $D_{\infty}$ is adequate but not fully abstract, and game semantics is fully abstract but requires machinery (arenas, strategies, innocence conditions) far beyond set theory. None of them eliminates the function space from the semantic domain; none avoids capture-avoiding substitution; and none dispenses with reduction rules or evaluation strategies.
Set-theoretic models of the $\lambda$-calculus
Several families of $\lambda$-calculus models avoid the inverse-limit construction of Scott’s $D_{\infty}$ by working in set-theoretic rather than domain-theoretic settings. Graph models (Engeler, 1981; Plotkin, 1993) take a web $A$ equipped with an injection $\iota:\mathcal{P}_{\mathrm{fin}}(A)\times A\hookrightarrow A$ and set $D=\mathcal{P}(A)$. Application is $f\cdot x=\{\,b\mid\exists\,S\subseteq x,\;\iota(S,b)\in f\,\}$: the function space $D\to D]$ is not constructed explicitly but encoded inside $D$ via $\iota$. The construction uses only powerset lattices, no Scott topology or continuity conditions. Graph models are sensible (they equate all unsolvable terms) but not fully abstract: Bucciarelli and Salibra (Bucciarelli and Salibra, 2008) proved that the greatest sensible graph theory equals the Böhm tree theory $\mathcal{B}$, which is strictly coarser than observational equivalence for the lazy $\lambda$-calculus. Filter models (Barendregt et al., 1983; Coppo and Dezani-Ciancaglini, 1978) interpret a term as the filter (upward-closed, intersection-closed set) of all intersection types derivable for it. The collection of all such filters forms an $\omega$-algebraic lattice. Barendregt, Coppo, and Dezani-Ciancaglini proved that type assignment is sound and complete with respect to the filter $\lambda$-model. Dezani-Ciancaglini et al. (Dezani-Ciancaglini et al., 1998) showed that the $\lambda$-theory of the BCD filter model coincides exactly with Böhm tree equality: two terms receive the same set of types if and only if they have the same Böhm tree. Filter models are more discriminating than graph models but are still not fully abstract for the lazy $\lambda$-calculus, and require the additional machinery of intersection type theory and algebraic lattices. Relational models (Ehrhard, 2005; Bucciarelli and Ehrhard, 2001) arise from the relational semantics of linear logic: the exponential modality !! is interpreted via finite multisets $\mathcal{M}_{\mathrm{fin}}(A)$, giving a co-Kleisli category whose morphisms are relations from $\mathcal{M}_{\mathrm{fin}}(A)$ to $B$. The multiset structure corresponds to non-idempotent intersection types (de Carvalho, 2018), which carry quantitative information (bounds on reduction length) that idempotent systems lack. Ehrhard (Ehrhard, 2005) showed that in the finiteness-space refinement of this model, standard fixed-point operators are not finitary, and a standard reflexive object for the pure $\lambda$-calculus cannot exist in this category. For the lazy $\lambda$-calculus, full abstraction has been achieved only through game semantics (Abramsky and Ong, 1993; Abramsky and McCusker, 1995); none of the three set-theoretic families attains it. All three families are semantic models of the $\lambda$-calculus: they interpret $\lambda$-terms in an external mathematical structure and validate $\beta$-reduction as an equation in that structure. None of them is a self-contained calculus with its own primitives into which the $\lambda$-calculus embeds. Moreover, all three identify terms more coarsely than observational equivalence—graph and filter models coincide with Böhm tree equality, while relational models cannot even host a standard reflexive object (Ehrhard, 2005)—and none provides a fully abstract semantics for the lazy $\lambda$-calculus without resorting to game-theoretic machinery.
@@
Amin et al. (Amin et al., 2016) formalized path-dependent types in the DOT calculus, the theoretical foundation of Scala’s type system. DOT paths navigate through objects to reach nested type members, providing a rich dependent type structure. DOT’s self variable in the object constructor $\nu(x:T)\,d$ binds $x$ to a single object; when multiple inheritance routes lead to the same enclosing scope, the resulting intersection type is unrealizable and Scala rejects it at compile time (Appendix B). DOT stratifies its syntax into terms and types, with types further divided into abstract type members bounded by $L..U$ and concrete definitions; each stratum requires its own subtyping rules. Whether these three roles could be unified under a single construct remains an open question in the DOT line of work. DOT includes the $\lambda$-calculus as a sub-language; removing it while retaining path-dependent types would yield a calculus closer to pure mixin inheritance (see Section 8).
@@
The NixOS module system (Dolstra et al., 2010; NixOS Contributors, 2024) composes modules by recursively merging nested attribute sets, not by linearization or method-level override, which is the same deep-merge semantics that distinguishes tree-level inheritance from method-level composition in prior mixin and trait calculi. When two modules define the same nested path, their subtrees are merged rather than overridden. The NixOS module system incorporates deep merge, recursive self-reference, deferred evaluation of modules, and a rich type-checking layer on top. The same mechanism powers NixOS system configuration, Home Manager (Helgesson, 2017), nix-darwin (Jordan, 2017), flake-parts (Hensing, 2022), and dozens of other ecosystems, collectively managing configurations of arbitrary complexity. The one aspect that appears not to have been fully addressed is the semantics of self-reference under multiple inheritance. When two independent modules that inherit from the same deferred module are composed, each import introduces its own declaration of the shared options; the module system’s mergeOptionDecls rejects duplicate declarations with a static error (Appendix B). The module system cannot express that both routes contribute equally to the same option, that is, the multi-path situation that recursive merging naturally gives rise to, which motivated the observational semantics of Section 3.
@@
Plugin systems, component frameworks, and dependency injection frameworks enable software extensibility through declarative inheritance. The symmetric, associative nature of inheritance (Section 3) explains why such systems work: components can be inherited in any order without changing behavior. These systems implement inheritance-calculus patterns ad hoc.
@@
Union file systems (UnionFS, OverlayFS, AUFS) layer multiple directory hierarchies to present a unified view. Their semantics resembles inheritance-calculus in that later layers override earlier layers and files from all layers remain accessible, but without a formal foundation these systems resort to ad-hoc choices: asymmetric inheritance, non-commutative conflict resolution for scalar files, and no late-binding or dynamic dispatch across layers. We implemented inheritance-calculus directly as a union file system (Yang, 2025), obtaining late-binding semantics and dynamic dispatch for file lookups with no additional mechanism beyond the three primitives. In package managers, build systems, and OS distributions, external toolchains transform configuration into file trees; here, the file system serves as configuration to itself, and the transformation is inheritance: a file-system-as-compiler in which the directory tree is both the source program and the compilation target.